Major league baseball pitchers throw a pitch called a changeup, which looks like a fastball leaving the pitcher's hand but has less speed than a typical fastball. How much more kinetic energy does a 92 mph fastball have than a 78 mph changeup? Express your answer as a percentage of the kinetic energy of the changeup.
Answer #1

percentage of the kinetic energy = 39.1 %

Explanation

V1 = 92 mph

V2 = 78 mph

percentage of kinetic energy = (V12 - V22) / V22

percentage of kinetic energy = (922 - 782) / 782

percentage of kinetic energy = 0.391

percentage of kinetic energy = 39.1 %
Answer #2

Step 1: Recall the formula for kinetic energy (KE):

KE=12mv2

Since mass (m) is the same for both pitches, we can compare their kinetic energies using only their velocities.

Step 2: Write the ratio of the fastball's KE to the changeup's KE:

KEfastballKEchangeup=vfastball2vchangeup2=(9278)2

(9278)2=(1.1795)21.39

Step 3: Find the percentage difference:

Percentage Increase=(1.391)×100%=39%

Answer is :

The 92 mph fastball has about 39% more kinetic energy than the 78 mph changeup.


answered by: Harshwardhan kunal